OpenStudy (callisto):

integration question \[\int\limits_{o}^{\pi/2}\frac{1}{1+sinx} dx =1\]

OpenStudy (callisto):

\[\int\limits_{o}^{\pi/2}\frac{1}{1+sinx} dx =1\]

OpenStudy (callisto):

I got \[[tanx-secx]_{0}^{\pi/2}\] But don't know how to evaluate it :|

OpenStudy (ash2326):

How did you evaluate the integral?

OpenStudy (callisto):

Sub. the number but tan (pi/2) is undefined... so is sec(pi/2)

OpenStudy (ash2326):

I mean how did you do the integration? I differentiated the last but not getting 1/(1+sin x)

OpenStudy (callisto):

\[\int\frac{1}{1+sinx}dx = \int\frac{1-sinx}{1-sin^2x}dx\]\[ = \int\frac{1-sinx}{cos^2x}dx = \int sec^2x - tanxsecxdx\]\[ =tanx - secx +C\]

OpenStudy (callisto):

In this case... not C...Sorry.

OpenStudy (callisto):

But when I sub it, as I said, it is undefined. And I haven't learnt improper integral yet

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{\pi}{2}} \frac{1}{1 + \sin x}dx = \left . \frac{2 \sin \left( \frac{x}{2} \right)}{\sin \left( \frac{x}{2} \right) + \cos \left( \frac{x}{2} \right)}\right|_0^{\frac{\pi}{2}}\]

OpenStudy (anonymous):

use t-method. t=tan(x/2) sinx=2t/(1-t^2) dx=(2/(1-t^2)) int (1/(1+sinx))dx =int (2/(1+2t+t^2))dt =int (2/(t+1)^2)dt =2 int(1/(t+1)^2)d(t+1) =2int(t+1)^(-2) d(t+1) =2 (-1)(t+1)^(-1) when x=pi/2 t=tan(pi/4)=1 when x=0 t=tan0=0 therefore, int (1/(1+sinx))dx =2 (-1)(1+1)^(-1) - 2 (-1)(0+1)^(-1) =-1+2 =1

OpenStudy (ash2326):

We'll use limits here We have \[[\tan x-\sec x]_{0} ^{\pi/2}\] Lower limit is just -1 \[\lim_{x \to \pi/2 } \frac{\sin x-1}{\cos x}-(-1)\] write \[x= \pi/2+ h\], we'll change limits here \(h\to 0\), so \[\lim_{h \to 0} \frac{ \cos h-1}{\sin h} -(-1)\] Can you proceed now?

OpenStudy (ash2326):

\[\tan x- \sec= \frac{\sin x-1}{\cos x}\]

OpenStudy (anonymous):

sorry, denominator of sinx and dt should be 1+t^2

OpenStudy (anonymous):

just these two is wrong. The other calculations are correct

OpenStudy (ash2326):

\[\lim_{h \to 0} \frac{-2\sin (h^2/2)}{2 \sin h/2 \cos h/2}=\lim_{h \to 0}\frac{-\sin h/2}{\cos h/2}= \frac{0}{1}=0\] so we get \[0-(-1)=1\]

OpenStudy (ash2326):

\[\cos h-1= -2\sin^2 h/2\] \[\sin h= 2\cos h/2 \times \sin h/2\]

OpenStudy (ash2326):

Did you get it? I have messed up a little

OpenStudy (ash2326):

Whenever during evaluation of limits, you are getting indefinite forms. Apply limits to evaluate that

OpenStudy (callisto):

I think I understand it... Thanks!

OpenStudy (ash2326):

Welcome:D